Sandy Bank, Inc., makes one model of wooden canoe. And, the information for it follows:

Number of canoes produced and sold 550 750 900
Total costs
Variable costs $ 110,000 $ 150,000 $ 180,000
Fixed costs $ 99,000 $ 99,000 $ 99,000
Total costs $ 209,000 $ 249,000 $ 279,000
Cost per unit
Variable cost per unit $ 200.00 $ 200.00 $ 200.00   
Fixed cost per unit 180.00 132.00 110.00   
Total cost per unit $ 380.00 $ 332.00 $ 310.00   


Required:
1.
Suppose that Sandy Bank raises its selling price to $500 per canoe. Calculate its new break-even point in units and in sales dollars. (Do not round intermediate calculations. Round your final answers to nearest whole number.)

New Break-Even Units
Break-Even Sales Revenue



2. If Sandy Bank sells 1,590 canoes, compute its margin of safety in dollars and as a percentage of sales. (Use the new sales price of $500.) (Round your answers to the nearest whole number.)

Margin of safety
Percentage of Sales



3. Calculate the number of canoes that Sandy Bank must sell at $500 each to generate $130,000 profit. (Round your answer to the nearest whole number.)

Target Sals Units
X X
X X
Answer #1

2) New break even point = 99000/(500-200) = 330 Units

Break even sales revenue = 330*500 = $165000

3) Margin of safety = (1590-330)*500 = $630000

Margin of safety percentage = 1590-330/1590 = 79%

4) Target sales unit = (130000+99000)/300 = 763 Units